OpenStudy (anonymous):

I need help in calc 2 please! Arc length

OpenStudy (anonymous):

https://books.google.com/books?id=PTELAAAAQBAJ&pg=PA532&lpg=PA532&dq=power+lines+are+constructed+by+stringing+wire+between+supports+and+ad&source=bl&ots=_wzBALOMoG&sig=liLxdakgthyG-dTK5Q1wi8-5-KU&hl=en&sa=X&ei=NcfwVOiFKsiqggTlyYOADQ&ved=0CCkQ6AEwAw#v=onepage&q=power%20lines%20are%20constructed%20by%20stringing%20wire%20between%20supports%20and%20ad&f=false Is the problem of the power lines dy/dx = (T/(ug)) sinh(ugx/T)(ug/T) = sinh(ugx/T) 1 + (dy/dx)^2 = 1 + sinh^2(ugx/T) = cosh^2(ugx/T) [1 + (dy/dx)^2]^(1/2)= cosh(ugx/T) integral of [1 + (dy/dx)^2]^(1/2) dx = integral of cosh(ugx/T) dx = [T/(ug)] sinh(ugx/T)

OpenStudy (anonymous):

is that good for the letter a?